OpenStudy (anonymous):

If a=5 and b=12 use the pythagorean theorem to solve for c

OpenStudy (anonymous):

13

OpenStudy (anonymous):

13

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 5^2 + 12^2 = c^2 25 + 144 = c^2 169 = c^2 c^2 = 169 c = sqrt(169) c = 13 So the solution is c =13

OpenStudy (anonymous):

what you need to do is : find a^2 + b^2 = c^2. therefore it will be 5^2 +12^2 ==169. squareroot that to get 13